OpenStudy (anonymous):

Hey! I can't remember the process to calculate this integral..., I have the integral table but I want to know the procedure. Thank! Integral of: sqrt(y^2 -9) dy

OpenStudy (anonymous):

try \(u=3\sin(\theta)\)

OpenStudy (loser66):

y =3 sec \(\theta\)

OpenStudy (anonymous):

Thanks! :)